OpenStudy (anonymous):

Trevor's dad decided to get him a cell phone for his birthday. Cell Plus has a plan that costs $20.00 per month plus an additional $0.06 per minute. Cell Best has a plan that costs $30.00 per month plus an additional $0.05 per minute. How many minutes can Trevor talk and both cell plans cost the same amount? 100 1000 500 50

OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 2x + y = 3 -3x + 4y = 1 (1, 1) (-1, -1) (1, -1) (-1, 1) i really need help on these 2 questions i really don't get how to do them if you could please help me!! that would be great!! thnx!! --Let the no of minutes be x 20 + 0.06x = 30 + 0.05x 0.01x = 10 x = 10 / 0.01 = 1000 minutes ANSWER CHECK 20 + 0.06*1000 = 80 $ 30 + 0.05*1000 = 80$ 2) 2x + y = 3 y = 3 -2x substitute in equation 2 -3x + 4 ( 3 - 2x) = 1 -3x +12 -8x = 1 11x = 11 x = 1 y = 1 ANSWER is A)

OpenStudy (anonymous):

its a.